OpenStudy (anonymous):

Solve for x -3 1/2 = 1/2x + 1/2x + x. -5 1/2 -1 3/4 -1 1/2

OpenStudy (anonymous):

\[\huge -3\frac12=\frac12\cdot x+\frac12 \cdot x+x\] is this the given?

OpenStudy (anonymous):

-7/2 x = x/2 + x/2 + x -7=x+x+2x(multiplying 2 on both sides) 4x=-7 x=-7/4=-1 3/4 third one...

OpenStudy (anonymous):

yes ,i think that is given...kreshnik

OpenStudy (anonymous):

@Sarkar how do you know there's \[\LARGE \frac12\cdot x\] instead of \[\LARGE \frac12x\] that wouldn't be : \[\LARGE \frac x2\] but \[\LARGE \frac12 +x \]

OpenStudy (anonymous):

@CoreyContini -please use the draw option to write the question properly....

OpenStudy (anonymous):

I guess @Sarkar 's solution is correct then... well done.